graph the relation x+lyl=3
Answer by lwsshak3(11628)

graph the relation x+lyl=3
rewrite equation:
|y|=-x+3
like all absolute value problems you must consider two conditions:
y=-x+3
-y=-x+3 or y=x-3
..
-x+3 is an equation of a straight line with slope=-1 and y-intercept=3
However, x<3 to keep y>0, so this line starts at (3,0) and continues to ∞ leftward up
..
The second condition, y=x-3, is a straight line with slope=1 and y-intercept=-3
Again, x<3 to keep y<0, so this line also starts at (3,0) and continues to ∞ leftward down
..
So what you have is a "V-shape" curve opening leftward with the apex at (3,0)
..
Sorry, I don't have the means to draw the graph for you. Hope this helps.